SIGNATURE OF CONFORMATIONAL RELAXATION IN THE MICROWAVE SPECTRUM OF ETHYL TRIFLUOROACETATE, (CF_{3}COO-CH_{2}CH_{3})

NANCY S. TRUE, Department of Chemistry, University of California, Davis, CA 95616; \underline{ROBERT K. BOHN}, Dept. of Chemistry, University of Connecticut, Storrs, CT 06269-3060.

Ethyl trifluoroacetate is a nearly prolate symmetric top, κ = -0.95, which displays three series of a-type R-branch bands under low resolution conditions in a Stark-modulated spectrometer.^{a} Since the molecule exists in only two stable conformers, the origin of the third band series has remained a puzzle for thirty years. We show that the third band series arises from rotational spectra of species undergoing exchange between the two stable forms. This is a purely microwave analog of conformational relaxation which has been observed in NMR spectra under different time, potential energy, and rate regimes.

aN. S. True, R. K. Bohn, J. Amer. Chem. Soc. 98, 1188 (1976).